# A delayed interior area-to-height estimate for the Curve Shortening Flow

### v1

@@ 5.3 Some comments on the work of Chou & Kwong
(5.43) $\|u-\widetilde{u}\|_{\mathrm{C}^{0}([a,b];\mathrm{L}^{p}(J_{r})\!)}=\sup_{t\in[a,b]}\|u(t)-\widetilde{u}(t)\|_{\mathrm{L}^{p}(J_{r})}\leq\varepsilon.\qed$
@@ 6 Towards a correspondence
In this section, we shall use this additional property, that each of the GCSFs generated by Theorem 1.13 is dominated by a positive GCSF, to further a correspondence between dominated GCSFs and $\mathcal{M}_{*}(\mathbb{R})$, following on from Chou & Kwong ([CK20, Appendix]). For the analogous correspondence in the RF setting, see [PT25, Theorem 1.7].
@@
    For each dominated GCSF $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$, there exists a unique $\nu\in\mathcal{M}_{*}(\mathbb{R})$ (so-called the initial trace in [CK20]) for which $u(t)\mathscr{L}^{1}\rightharpoonup\nu\in\mathcal{M}_{*}(\mathbb{R})$ as $t\searrow 0$.
Proof.
@@
(6.4) $\left|\,\int_{x\in\mathbb{R}}\varphi(x)u(x,t)\,\mathrm{d}x-\int_{x\in\mathbb{R}}\varphi(x)u(x,s)\,\mathrm{d}x\,\right|\leq C(\varphi)(t-s),$
@@
(6.5) $\varphi\mapsto L\varphi=\lim_{t\searrow 0}\int_{x\in\mathbb{R}}\varphi(x)u(x,t)\,\mathrm{d}x:\mathrm{C}^{\infty}_{\mathrm{c}}(\mathbb{R})\mapsto\mathbb{R}$
is well-defined. The map $L$ also inherits the linearity of the integral in the limit.
@@
(6.6) $\left|\,\int_{x\in\mathbb{R}}\varphi(x)u(x,t)\,\mathrm{d}x\,\right|$ $\leq\int_{x\in\mathbb{R}}\|\varphi\|_{\mathrm{C}^{0}(\mathcal{K})}\chi\cdot U(x,t)\,\mathrm{d}x=\|\varphi\|_{\mathrm{C}^{0}(\mathcal{K})}\int_{x\in\mathbb{R}}\chi(x)U(x,t)\,\mathrm{d}x$
@@
(6.7) $|L\varphi|\leq M_{\mathcal{K}}\|\varphi\|_{\mathrm{C}^{0}(\mathcal{K})}.$
The Riesz Representation Theorem ([EG15, Theorem 1.38]) ensures that $L$ is given by integration against a real–valued Radon measure $\nu\in\mathcal{M}(\mathbb{R})$. To show that $\nu\in\mathcal{M}_{*}(\mathbb{R})$ (i.e. that $\nu$ is non-atomic), it suffices to show for all $p\in\mathbb{R}$ and $\varepsilon>0$ that $|\nu(V)|\leq\varepsilon$ for all sufficiently small and open $V\ni p$. Given a bump function $\varphi\in\mathrm{C}^{\infty}_{\mathrm{c}}(\mathbb{R};[0,1])$ around $p$ with $C(\varphi)=\pi$, note that
(6.8) $\left|\,\int_{x\in\mathbb{R}}\varphi(x)\,\mathrm{d}\nu(x)\right|\leq\int_{x\in\mathbb{R}}\varphi(x)U\left(x,\frac{\varepsilon}{2\pi}\right)\,\mathrm{d}x+\frac{\varepsilon}{2}.$
If we choose $R_{p}>0$ so small that
(6.9) $\int_{x\in J_{2R_{p}}(p)}U\left(x,\frac{\varepsilon}{2}\right)\,\mathrm{d}x\leq\frac{\varepsilon}{2},$
@@
(6.10) $\left|\,\int_{x\in\mathbb{R}}\varphi(x)\,\mathrm{d}\nu(x)\right|\leq\int_{J_{2R_{p}}(p)}U\left(\frac{\varepsilon}{2\pi}\right)\,\mathrm{d}\mathscr{L}^{1}+\frac{\varepsilon}{2}\leq\varepsilon$
whenever $\varphi$ is supported in $J_{2R_{p}}(p)$. In particular, given any open $V\subset J_{R_{p}}(p)$ containing $p$, if $(\chi^{j})_{j\in\mathbb{N}}\subset\mathrm{C}^{\infty}_{\mathrm{c}}(J_{2R_{p}}(p);[0,1])$ is such that $\chi^{j}\to\chi_{V}$ point-wise as $j\nearrow\infty$, then by the Dominated Convergence Theorem one can pass to the limit in (6.10) to deduce the required
(6.11) $|\nu(V)|=\lim_{j\nearrow\infty}\left|\,\int_{x\in\mathbb{R}}\chi^{j}(x)\,\mathrm{d}\nu(x)\right|\leq\varepsilon.\qed$
@@
(6.12) $\lim_{t\searrow 0}\int_{x\in\mathbb{R}}\varphi(x)u(x,t)\,\mathrm{d}x$
in the case $\varphi\in\mathrm{C}^{1}_{\mathrm{c}}(\mathbb{R})$ can be understood as the distributional derivative of a unique $\mathrm{L}^{\infty}_{\mathrm{loc}}(\mathbb{R})$–function. In the positive case $u\geq 0$ on ×$\mathbb{R}\times(0,T)$, they recovered the characterisation as an element of $\mathcal{M}_{*}(\mathbb{R})$ given here.
@@
Appendix A Angenent’s Intersection Principle
@@
Let $\gamma^{1,2}:[0,1]\times[0,T)\mapsto\mathbb{R}^{2}$ be two distinct CSFs for which
(A.2) $\partial\gamma^{1}(t)\cap\gamma^{2}(t)=\gamma^{1}(t)\cap\partial\gamma^{2}(t)=\emptyset$
@@
We say that a CSF $\gamma:J\times[0,T)\mapsto\mathbb{R}^{2}$ is uniformly proper if every restriction $\gamma:J\times[0,t]\mapsto\mathbb{R}^{2}$ for $0<t<T$ is a proper map.
Peachey’s subclass of CSFs possesses the important feature that, Ecker–Huisken flows of locally Lipschitz functions and Chou–Zhu flows of curves whose ends are graphical are automatically uniformly proper—see [Sob24, Remark 2.2.31 and Corollary 2.4.15].
@@
We say that two CSFs $\gamma^{1,2}:J^{1,2}\times[0,T)\mapsto\mathbb{R}^{2}$ have disjoint ends if (A.2) holds and if there exists a compact set $\mathcal{K}\subset\mathbb{R}^{2}$ for which
(A.5) $\gamma^{1}(t)\cap\gamma^{2}(t)\cap(\mathbb{R}^{2}\setminus\mathcal{K})=\emptyset$
@@
Armed with Definitions A.3 and A.4, we have the following version of Angenent’s Intersection Principle, which shall be the main version used in this paper.
@@
Let $\gamma^{1,2}:J^{1,2}\times[0,T)\mapsto\mathbb{R}^{2}$ be two distinct uniformly proper CSFs, starting from the locally Lipschitz curves $\gamma^{1,2}_{0}$, which have disjoint ends. Then the number of intersections between $\gamma^{1}(t)$ and $\gamma^{2}(t)$ is finite function of $t\in(0,T)$, and a decreasing function of $t\in[0,T)$ which strictly decreases whenever $\gamma^{1}(t)$ and $\gamma^{2}(t)$ meet tangentially. Moreover, suppose that $\gamma^{1}_{0}$ and $\gamma^{2}_{0}$ intersect one-sidedly [^6] An intersection $q$ of $\gamma^{1}_{0}$ and $\gamma^{2}_{0}$, with $\gamma^{1}_{0}$ of class $\mathrm{C}^{1}$ near $q$, is one-sided if, in a neighbourhood of $q$ we can write sub-arcs of $\gamma^{1,2}_{0}$ as ordered graphs over a common smooth arc with the ordering becoming strict towards the end-points. $m$ times and otherwise intersect $n$ times, within regions where at least one of the curves is $\mathrm{C}^{1}$; then the number of intersections for strictly positive times is at most $n$.
@@
The uniformly proper condition and the assumption of disjoint ends allows one to reduce to the case that $J^{1,2}$ are both compact, at which point the classical Angenent’s Intersection Principle (Theorem A.1) applies. The Strong Maximum Principle implies that one-sided initial intersections disappear instantaneously. If there are infinitely many intersections which are not one-sided, then the final assertion is trivial. For finitely many initial intersections that are not one-sided, a closer look at Angenent’s [Ang88, Lemmas 5.3 and 5.5] reveals that each one of these must not bifurcate along the flow for a short length of time. For details, see [Sob24, Theorem 2.4.10]. ∎
An immediate consequence of Angenent’s Intersection Principle is the following version of the Avoidance Principle.
@@
Let $\gamma^{1,2}:J^{1,2}\times[0,T)\mapsto\mathbb{R}^{2}$ be two uniformly proper CSFs, starting from the locally Lipschitz curves $\gamma^{1,2}_{0}$, which have disjoint ends. Suppose that $\gamma^{1}_{0}$ and $\gamma^{2}_{0}$ only intersect one-sidedly within regions where at least one of the curves is $\mathrm{C}^{1}$; then for strictly positive times $t>0$ the curves $\gamma^{1}(t)$ and $\gamma^{2}(t)$ are disjoint.
Appendix B Real–valued Radon measures
@@
By a real–valued Radon measure ([EG15, §1.8]) on $\mathbb{R}$, we mean a Radon measure $\mu$ on $\mathbb{R}$ paired with an $\mu$–measurable function $\sigma:\mathbb{R}\mapsto\left\{\pm 1\right\}$. We write $\nu=\sigma\mu$ and $|\nu|=\mu$. We denote the space of real–valued Radon measures by $\mathcal{M}(\mathbb{R})$ and the space of Radon measures by $\mathcal{M}^{+}(\mathbb{R})$.
@@
(B.1) $\int_{\mathbb{R}}\varphi\,\mathrm{d}\nu_{n}\to\int_{\mathbb{R}}\varphi\,\mathrm{d}\nu$
@@
We say that a real–valued Radon measure $\nu$ is non-atomic if $\nu(\left\{x\right\})=0$ for all $x\in\mathbb{R}$. (Note that Chou & Kwong call such a measure continuous in [CK20].) We denote the space of non-atomic real–valued Radon measures by $\mathcal{M}_{*}(\mathbb{R})$, and the space of non-atomic Radon measures by $\mathcal{M}^{+}_{*}(\mathbb{R})$. Clearly $\nu\in\mathcal{M}_{*}(\mathbb{R})$ if and only if $|\nu|\in\mathcal{M}^{+}_{*}(\mathbb{R})$.
Appendix C Evans & Spruck’s interior gradient estimate
@@
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{2R}\times[0,2T])$ be a GCSF for which $\|u\|_{\mathrm{L}^{\infty}(J_{2R}\times[0,2T])}\leq L$. Then, for any $0<\sigma<T$, there exists a constant $C=C(R,\sigma,T,L)<\infty$ such that
(C.2) $\|\partial_{x}u\|_{\mathrm{L}^{\infty}(J_{R}\times[\sigma,T])}\leq C\left(r,\sigma,T,\|u\|_{\mathrm{L}^{\infty}(J_{2R}\times[0,2T])}\right).$
Sketch proof.
@@
(C.3) $|\partial_{x}u(0,\tau)|\leq C\left(R,\tau,\|u\|_{\mathrm{L}^{\infty}(J_{R}\times[0,2\tau])}\right).$
To obtain (C.2), we apply (C.3) to $J_{R}\times[0,2s]\ni(x,t)\mapsto\widetilde{u}(x,t)=u(y+x,t)\in\mathbb{R}$ for every $(y,s)\in J_{R}\times[\sigma,T]$. ∎

### v2

@@ 5.3 Some comments on the work of Chou & Kwong
$\|u-\widetilde{u}\|_{\mathrm{C}^{0}([a,b];\mathrm{L}^{p}(J_{r})\!)}=\sup_{t\in[a,b]}\|u(t)-\widetilde{u}(t)\|_{\mathrm{L}^{p}(J_{r})}\leq\varepsilon.\qed$ (5.42)
Remark 5.43.
While Theorem 5.38, together with [CK20, Theorem A], fully resolves the question of well-posed-ness of the GCSF from $\mathrm{L}^{p}_{\mathrm{loc}}$–initial data when $p>1$ is strictly greater than $1$, the question of uniqueness and continuous dependence from $\mathrm{L}^{1}_{\mathrm{loc}}$–initial data, i.e. the case when $p=1$, is still open. Since Proposition 5.25 in the case $p=1$ is sharp—see Remark 5.27—the author expects a proof of uniqueness from $\mathrm{L}^{1}_{\mathrm{loc}}$–initial data to require a novel estimate.
@@ 6 Towards a correspondence
In this section, we shall use this additional property, that each of the GCSFs generated by Theorem 1.14 is dominated by a positive GCSF, to further a correspondence between dominated GCSFs and $\mathcal{M}_{*}(\mathbb{R})$, following on from Chou & Kwong in [CK20, Appendix]. For the analogous correspondence in the RF setting, see [PT25, Theorem 1.7].
@@
    For each dominated GCSF $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$, there exists a unique $\nu\in\mathcal{M}_{*}(\mathbb{R})$, so-called the initial trace in [CK20], for which $u(t)\mathscr{L}^{1}\rightharpoonup\nu\in\mathcal{M}_{*}(\mathbb{R})$ as $t\searrow 0$.
?proofname?.
@@
$\left|\,\int_{x\in\mathbb{R}}\varphi(x)u(x,t)\,\mathrm{d}x-\int_{x\in\mathbb{R}}\varphi(x)u(x,s)\,\mathrm{d}x\,\right|\leq C(\varphi)(t-s),$ (6.4)
@@
$\varphi\mapsto L\varphi=\lim_{t\searrow 0}\int_{x\in\mathbb{R}}\varphi(x)u(x,t)\,\mathrm{d}x:\mathrm{C}^{\infty}_{\mathrm{c}}(\mathbb{R})\mapsto\mathbb{R}$ (6.5)
is well-defined, and is entirely determined by the flow $u$. The map $L$ inherits the linearity of the integral in the limit.
@@
$\left|\,\int_{x\in\mathbb{R}}\varphi(x)u(x,t)\,\mathrm{d}x\,\right|$ $\leq\int_{x\in\mathbb{R}}\|\varphi\|_{\mathrm{C}^{0}(\mathcal{K})}\chi\cdot U(x,t)\,\mathrm{d}x=\|\varphi\|_{\mathrm{C}^{0}(\mathcal{K})}\int_{x\in\mathbb{R}}\chi(x)U(x,t)\,\mathrm{d}x$ (6.6)
@@
$|L\varphi|\leq M_{\mathcal{K}}\|\varphi\|_{\mathrm{C}^{0}(\mathcal{K})}.$ (6.7)
The Riesz Representation Theorem ([EG15, Theorem 1.38]) ensures that $L$ is given by integration against a unique real–valued Radon measure $\nu\in\mathcal{M}(\mathbb{R})$. To show that $\nu\in\mathcal{M}_{*}(\mathbb{R})$, i.e. that $\nu$ is non-atomic, it suffices (by outer regularity) to show for all $z\in\mathbb{R}$ and $\varepsilon>0$ that $|\nu(V)|\leq\varepsilon$ for all sufficiently small and open $V\ni z$. Given a bump function $\varphi\in\mathrm{C}^{\infty}_{\mathrm{c}}(\mathbb{R};[0,1])$ around $z$ with $C(\varphi)=\pi$, note that
$\left|\,\int_{x\in\mathbb{R}}\varphi(x)\,\mathrm{d}\nu(x)\right|\leq\int_{x\in\mathbb{R}}\varphi(x)U\left(x,\frac{\varepsilon}{2\pi}\right)\,\mathrm{d}x+\frac{\varepsilon}{2}.$ (6.8)
If we choose $R_{z}>0$ so small that
$\int_{x\in J_{2R_{z}}(z)}U\left(x,\frac{\varepsilon}{2\pi}\right)\,\mathrm{d}x\leq\frac{\varepsilon}{2},$ (6.9)
@@
$\left|\,\int_{x\in\mathbb{R}}\varphi(x)\,\mathrm{d}\nu(x)\right|\leq\int_{x\in J_{2R_{z}}(z)}U\left(x,\frac{\varepsilon}{2\pi}\right)\,\mathrm{d}x+\frac{\varepsilon}{2}\leq\varepsilon$ (6.10)
whenever $\varphi$ is supported in $J_{2R_{z}}(z)$. In particular, given any open $V\subset J_{R_{z}}(z)$ containing $z$, if $(\chi^{j})_{j\in\mathbb{N}}\subset\mathrm{C}^{\infty}_{\mathrm{c}}(J_{2R_{z}}(z);[0,1])$ is such that $\chi^{j}\to\chi_{V}$ pointwise as $j\nearrow\infty$, then by the Dominated Convergence Theorem one can pass to the limit in (6.10) to deduce the required
$|\nu(V)|=\lim_{j\nearrow\infty}\left|\,\int_{x\in\mathbb{R}}\chi^{j}(x)\,\mathrm{d}\nu(x)\right|\leq\varepsilon.\qed$ (6.11)
@@
$\lim_{t\searrow 0}\int_{x\in\mathbb{R}}\varphi(x)u(x,t)\,\mathrm{d}x$ (6.12)
in the case $\varphi\in\mathrm{C}^{1}_{\mathrm{c}}(\mathbb{R})$ can be understood as the distributional derivative of a unique $\mathrm{L}^{\infty}_{\mathrm{loc}}(\mathbb{R})$–function. And, in the positive case $u\geq 0$, they similarly characterised the limit as an element of $\mathcal{M}_{*}(\mathbb{R})$.
@@
?appendixname? A Angenent’s Intersection Principle
@@
Let $0,T)\ni t\mapsto\gamma^{1}(t),\gamma^{2}(t)\subset\mathbb{R}^{2}$ be two distinct CSFs, each defined over either the circle $\mathbb{S}^{1}$ or the closed interval $[0,1]$, for which
$\partial\gamma^{1}(t)\cap\gamma^{2}(t)=\gamma^{1}(t)\cap\partial\gamma^{2}(t)=\emptyset$ (A.2)
@@
We say that a CSF $\gamma:\Omega\times[0,T)\mapsto\mathbb{R}^{2}$ is uniformly proper if every restriction $\gamma:\Omega\times[0,t]\mapsto\mathbb{R}^{2}$ for $0<t<T$ is a proper map.
Peachey’s subclass of CSFs possesses the important feature that, Ecker–Huisken flows of locally Lipschitz functions and Chou–Zhu flows of curves whose ends are graphical are automatically uniformly proper—see [Sob24, Remark 2.2.31 and Corollary 2.4.15]. Clearly compact CSFs are automatically uniformly proper.
@@
We say that two CSFs $0,T)\ni t\mapsto\gamma^{1}(t),\gamma^{2}(t)\subset\mathbb{R}^{2}$ have disjoint ends if (A.2) holds and if there exists a compact set $\mathcal{K}\Subset\mathbb{R}^{2}$ for which
$\gamma^{1}(t)\cap\gamma^{2}(t)\cap(\mathbb{R}^{2}\setminus\mathcal{K})=\emptyset$ (A.5)
@@
Armed with Definitions A.3 and A.4, we have the following version of Angenent’s Intersection Principle, which shall be the main version used in this paper:
@@
Let $0,T)\ni t\mapsto\gamma^{1}(t),\gamma^{2}(t)\subset\mathbb{R}^{2}$ be two distinct uniformly proper CSFs, starting from the locally Lipschitz curves $\gamma^{1}_{0}$, $\gamma^{2}_{0}$ respectively, which have disjoint ends. Then the number of intersections between $\gamma^{1}(t)$ and $\gamma^{2}(t)$ is finite function of $t\in(0,T)$, and a decreasing function of $t\in[0,T)$ which strictly decreases whenever $\gamma^{1}(t)$ and $\gamma^{2}(t)$ meet tangentially. Moreover, suppose that $\gamma^{1}_{0}$ and $\gamma^{2}_{0}$ intersect one-sidedly [^4] An intersection $q$ of $\gamma^{1}_{0}$ and $\gamma^{2}_{0}$, with $\gamma^{1}_{0}$ of class $\mathrm{C}^{1}$ near $q$, is one-sided if, in a neighbourhood of $q$ we can write sub-arcs of $\gamma^{1}_{0}$ and $\gamma^{2}_{0}$ as ordered graphs over a common smooth arc with the ordering becoming strict towards the end-points. $m$ times and otherwise intersect $n$ times, within regions where at least one of the curves is $\mathrm{C}^{1}$; then the number of intersections for strictly positive times is at most $n$.
@@
The uniformly proper condition and the assumption of disjoint ends allows one to reduce to the case that $\gamma^{1}(t)$ and $\gamma^{2}(t)$ are both compact for each $t\in[0,T)$, at which point the classical Angenent’s Intersection Principle (Theorem A.1) applies. The Strong Maximum Principle implies that one-sided initial intersections disappear instantaneously. If there are infinitely many intersections which are not one-sided, then the final assertion is trivial. For finitely many initial intersections that are not one-sided, a closer look at Angenent’s [Ang88, Lemmas 5.3 and 5.5] reveals that each one of these must not bifurcate along the flow for a short length of time. For details, see [Sob24, Theorem 2.4.10]. ∎
An immediate consequence of Angenent’s Intersection Principle is the following version of the Avoidance Principle:
@@
Let $0,T)\ni t\mapsto\gamma^{1}(t),\gamma^{2}(t)\subset\mathbb{R}^{2}$ be two uniformly proper CSFs, starting from the locally Lipschitz curves $\gamma^{1}_{0}$, $\gamma^{2}_{0}$ respectively, which have disjoint ends. Suppose that $\gamma^{1}_{0}$ and $\gamma^{2}_{0}$ only intersect one-sidedly within regions where at least one of the curves is $\mathrm{C}^{1}$. Then for strictly positive times $t>0$ the curves $\gamma^{1}(t)$ and $\gamma^{2}(t)$ are disjoint.
?appendixname? B Real–valued Radon measures
@@
By a real–valued Radon measure ([EG15, §1.8]) on $\mathbb{R}$, we mean a Radon measure $\mu$ on $\mathbb{R}$ paired with a $\mu$–measurable function $\sigma:\mathbb{R}\mapsto\left\{\pm 1\right\}$. We write $\nu=\sigma\mu$ and $|\nu|=\mu$. We denote the space of real–valued Radon measures by $\mathcal{M}(\mathbb{R})$ and the space of Radon measures by $\mathcal{M}^{+}(\mathbb{R})$.
@@
$\int_{\mathbb{R}}\varphi\,\mathrm{d}\nu_{n}\to\int_{\mathbb{R}}\varphi\,\mathrm{d}\nu$ (B.1)
@@
We say that a real–valued Radon measure $\nu$ is non-atomic if $\nu(\left\{x\right\})=0$ for all $x\in\mathbb{R}$; note that Chou & Kwong call such a measure continuous in [CK20]. We denote the space of non-atomic real–valued Radon measures by $\mathcal{M}_{*}(\mathbb{R})$, and the space of non-atomic Radon measures by $\mathcal{M}^{+}_{*}(\mathbb{R})$. Clearly $\nu\in\mathcal{M}_{*}(\mathbb{R})$ if and only if $|\nu|\in\mathcal{M}^{+}_{*}(\mathbb{R})$.
?appendixname? C Evans & Spruck’s interior gradient estimate
@@
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{2R}\times[0,2T])$ be a GCSF for which $\|u\|_{\mathrm{L}^{\infty}(J_{2R}\times[0,2T])}\leq L$. Then, for any $0<S<T$, there exists a constant $C(R,S,T,L)<\infty$ such that
$\|\partial_{x}u\|_{\mathrm{L}^{\infty}(J_{R}\times[S,T])}\leq C(R,S,T,L).$ (C.2)
?proofname?.
@@
$|\partial_{x}u(0,s)|\leq C\left(R,s,\|u\|_{\mathrm{L}^{\infty}(J_{R}\times[0,2s])}\right).$ (C.3)
To obtain (C.2), we apply (C.3) to $J_{R}\times[0,2s]\ni(x,t)\mapsto\widetilde{u}(x,t)=u(y+x,t)\in\mathbb{R}$ for every $(y,s)\in J_{R}\times[S,T]$. ∎
?refname?
  • [Ang88] S. B. Angenent, The zero set of a solution of a parabolic equation, Journal für die Reine und Angewandte Mathematik (Crelles Journal) 1988 (1988), no. 390, 79–96.
  • [Ang90]   , Parabolic equations for curves on surfaces: Part I. Curves with $p$-integrable curvature, Annals of Mathematics 132 (1990), no. 3, 451–483.
  • [Ang91]   , Parabolic equations for curves on surfaces: Part II. Intersections, blow-up and generalized solutions, Annals of Mathematics 133 (1991), no. 1, 171–215.
  • [BR24] T. Bourni and M. Reiris, Compact curve shortening flow solutions out of non-compact curves, International Mathematics Research Notices 2024 (2024), no. 1, 804–816.
  • [CK01] K.-S. Chou and Y.-C. Kwong, On quasilinear parabolic equations which admit global solutions for initial data with unrestricted growth, Calculus of Variations and Partial Differential Equations 12 (2001), 281–315.
  • [CK20]   , General initial data for a class of parabolic equations including the curve shortening problem, Discrete and Continuous Dynamical Systems 40 (2020), no. 5, 2963–2986.
  • [Clu04] J. Clutterbuck, Parabolic equations with continuous initial data, Ph.D. thesis, Australian National University, 2004.
  • [CM04] T. H. Colding and W. P. Minicozzi II, Sharp estimates for mean curvature flow of graphs, Journal für die Reine und Angewandte Mathematik (Crelles Journal) 2004 (2004), no. 574, 187–195.
  • [CZ98] K.-S. Chou and X.-P. Zhu, Shortening complete plane curves, Journal of Differential Geometry 50 (1998), no. 3, 471–504.
  • [DS23] P. Daskalopoulos and M. Saez, Uniqueness of entire graphs evolving by mean curvature flow, Journal für die Reine und Angewandte Mathematik (Crelles Journal) 2023 (2023), no. 796, 201–227.
  • [EG15] L. C. Evans and R. F. Gariepy, Measure theory and fine properties of functions, revised 1 ed., Textbooks in Mathematics, Chapman & Hall/CRC, 2015.
  • [EH89] K. Ecker and G. Huisken, Mean curvature evolution of entire graphs, Annals of Mathematics 130 (1989), no. 3, 453–471.
  • [EH91]   , Interior estimates for hypersurfaces moving by mean curvarture, Inventiones Mathematicae 105 (1991), no. 1, 547–569.
  • [ES92] L. C. Evans and J. Spruck, Motion of level sets by mean curvature. III, The Journal of Geometric Analysis 2 (1992), no. 2, 121–150.
  • [GH86] M. E. Gage and R. S. Hamilton, The heat equation shrinking convex plane curves, Journal of Differential Geometry 23 (1986), no. 1, 69–96.
  • [Gra87] M. A. Grayson, The heat equation shrinks embedded plane curves to round points, Journal of Differential geometry 26 (1987), no. 2, 285–314.
  • [Gra89]   , Shortening embedded curves, Annals of Mathematics 129 (1989), no. 1, 71–111.
  • [Ham93] R. S. Hamilton, The Harnack estimate for the Ricci flow, Journal of Differential Geometry 37 (1993), no. 1, 225–243.
  • [Hui84] G. Huisken, Flow by mean curvature of convex surfaces into spheres, Journal of Differential Geometry 20 (1984), no. 1, 237–266.
  • [Ilm92] T. Ilmanen, Generalized flow of sets by mean curvature on a manifold, Indiana University Mathematics Journal 41 (1992), no. 3, 671–705.
  • [Kry96] N. V. Krylov, Lectures on elliptic and parabolic equations in Hölder spaces, Graduate Studies in Mathematics, vol. 12, American Mathematical Society, 1996.
  • [Lau13] J. Lauer, A new length estimate for curve shortening flow and low regularity initial data, Geometric and Functional Analysis 23 (2013), no. 6, 1934–1961.
  • [LSU88] O. A. Ladyženskaja, V. A. Solonnikov, and N. N. Ural’ceva, Linear and quasi-linear equations of parabolic type, Translations of Mathematical Monographs, vol. 23, American Mathematical Society, 1988.
  • [NT05] Y. Nagase and Y. Tonegawa, Interior gradient estimate for 1-D anisotropic curvature flow, Boletim da Sociedade Paranaense de Matemática 23 (2005), no. 1–2, 93–98.
  • [Pea22] L. T. Peachey, Geometric flows without boundary data at infinity, Ph.D. thesis, University of Warwick, 2022.
  • [PT25] L. T. Peachey and P. M. Topping, All two-dimensional expanding Ricci solitons, Journal of the London Mathematical Society 111 (2025), no. 3, e70072.
  • [Rud91] W. Rudin, Functional analysis, 2 ed., McGraw–Hill, 1991.
  • [Sob24] A. Sobnack, Geometric regularity properties of the Curve Shortening Flow, Ph.D. thesis, University of Warwick, 2024.
  • [ST24a] A. Sobnack and P. M. Topping, Delayed parabolic regularity for curve shortening flow, arXiv preprint arXiv:2408.04049 (2024).
  • [ST24b]   , Monotonicity of the modulus under curve shortening flow, arXiv preprint arXiv:2409.03098 (2024).
  • [ST26]   , The Harnack inequality without convexity for curve shortening flow, arXiv preprint arXiv:2601.13767 (2026).
  • [TY17] P. M. Topping and H. Yin, Sharp decay estimates for the logarithmic fast diffusion equation and the Ricci flow on surfaces, Annals of Partial Differential Equations 3 (2017), no. 6, 1–16.
  • [TY24a]   , Smoothing a measure on a Riemann surface using Ricci flow, Ars Inveniendi Analytica 2024 (2024), no. 4, DOI 10.15781/4x5c–9q97.
  • [TY24b]   , Uniqueness of Ricci flows from non-atomic Radon measures on Riemann surfaces, Proceedings of the London Mathematical Society 128 (2024), no. 6, e12600.